How do I do solve the problem 2x+y=9? I need to find 4 ordered pair solutions for this equation.
Answer by Alan3354(69443)

How do I do solve the problem 2x+y=9?
------
It's not a problem, it's an equation.
Solve for what?
You can solve for x:
2x+y=9
2x = -y + 9
x = -y/2 + 9/2
==============
You can solve for y:
y = -2x + 9
==============
I need to find 4 ordered pair solutions for this equation.
Use y = -2x + 9
Pick 4 values for x and find y.